OpenStudy (fanduekisses):

How to evaluate log of (1/12) base 2?

OpenStudy (fanduekisses):

\[\log _{2}\frac{ 1 }{128}\]

OpenStudy (fanduekisses):

\[2^{x}=\frac{ 1 }{128}\]

ganeshie8 (ganeshie8):

thats a very good start !

ganeshie8 (ganeshie8):

use below : 128 = 2^7

OpenStudy (fanduekisses):

so x=-7?

ganeshie8 (ganeshie8):

Correct !

OpenStudy (fanduekisses):

yay :D thanks ^_^

ganeshie8 (ganeshie8):

good job !!